# Cody D. Schimming, C. J. O. Reichhardt, and C. Reichhardt

### v1

@@ 3 Interaction between two active nematic defects
Here we analyze the interaction between active nematic defects with the simplifying assumption that only two defects exist in an infinite domain. In each case we set our coordinates so that the defect of interest is at polar coordinates $(r,\theta)$ and the other defect is at $(0,0)$. We will also assume “optimal” orientation between defects 44. This set of orientations minimizes the elastic free energy of the configuration. Further, given a global phase $\varphi_{0}$, the orientation of each defect may be determined by the defect positions:
@@ 3.1 Two $+1/2$ Defects
where $\mathbf{\hat{r}}$ is the unit vector pointing to the defect of interest from the other defect, $\mathbf{\hat{a}}=\left[\cos(2\theta-2\varphi_{0}),\sin(2\theta-2\varphi_{0})\right]$, and $\mathbf{\hat{b}}=\left[\cos(3\theta-4\varphi_{0}),\sin(3\theta-4\varphi_{0})\right]$. Components of Eq. (22) proportional to $\beta_{i}$ result from advection while components proportional to $\chi_{i}$ result from the interaction with shear flows. We note that the other defect will have exactly the opposite velocity of the defect under consideration.
@@
$v_{||}^{+}=\frac{1}{r}-\frac{\alpha}{16\ell_{h}^{2}}\chi_{1}\\ +\frac{\alpha}{4\ell_{h}}\left(-\pi+\beta_{1}+\beta_{2}\right)\cos(\theta-2\varphi_{0})-\frac{\alpha}{16\ell_{h}^{2}}\chi_{2}\cos(2\theta-4\varphi_{0})$ (23)
$v_{\perp}^{+}=-\frac{\alpha}{4\ell_{h}}\left(-\pi+\beta_{1}+\beta_{2}\right)\sin(\theta-2\varphi_{0})\\ +\frac{\alpha}{16\ell_{h}^{2}}\chi_{2}\sin(2\theta-4\varphi_{0})$ (24)
where we have suppressed the arguments of $\beta_{1},\beta_{2},\chi_{1},\chi_{2}$ for brevity. We first note that if the configuration is rotated by $\theta$, the angular arguments $\theta-2\varphi_{0}$ remain unchanged since under rotation by $\theta$, $\varphi_{0}\to\varphi_{0}+(1/2)\theta$. Thus, the decomposition predicts that two $+1/2$ defects will rotate around one another if $\theta\neq 2\varphi_{0}+n\pi$, and we may analyze the motion by rotating our axes so that $\theta=0$, with $\varphi_{0}$ determining the motion.
@@
The behavior of $r_{0}$ may be attributed primarily to the self-advection part of Eq. (23), which is the dominant term in the equation. It does not depend on the distance between defects, and it contributes most strongly at high activities and smaller $\varphi_{0}$. However, we note that the shear flow interaction contributes most strongly to the behavior near $\varphi_{0}=\pi/4$ where the advection terms do not contribute to Eq. (23). If $\lambda=0$, $r_{0}\to\infty$ for $\varphi_{0}=\pi/4$.
@@
The effective attraction between positive defects caused by hydrodynamic effects can lead to dynamical states in which $+1/2$ defects continuously rotate around one another, or even merge to form a stable higher order $+1$ defect, similar to behaviors predicted with a different analytical model in Ref. 45. While our calculations above predict this attraction for the simplest case of a system with only two defects, this behavior should, in principle, occur even when other defects are present in the system if the two defects are far from the other defects. Indeed, bound states of two $+1/2$ defects may be observed in continuum simulations at high activities, as illustrated in Fig. 4(a) where we show a time snapshot of the scalar order parameter and director field in a system with periodic boundary conditions, $\alpha/\ell_{h}^{2}=2$, and $\ell_{h}=10$ (for details of the continuum simulations, see the Appendix). We have marked the bound states of $+1/2$ defects in green boxes in Fig. 4(a), and for one of these boxes the subpanels provide several time snapshots illustrating two bound $+1/2$ defects rotating around one another and eventually merging. We note that the bound states are short-lived in the continuum simulation due to interactions with other defects. However, it has been shown in numerical simulations that such bound states may be longer lived and ordered if subjected to confinement 46, 47, 48. Partially stable bound $+1/2$ defect pairs have also been observed experimentally under circular confinement 49 and in recently discovered acoustically actuated active nematics with no confinement, where the activity may be much larger than in biologically based active nematics 50.
@@
Fig.  4: (a) Time snapshot of an active nematic continuum simulation with periodic boundary conditions, $\alpha/\ell_{h}^{2}=2$, and $\ell_{h}=10$. The color represents the strength of the scalar order parameter $S$ and the white lines indicate the local director. The green boxes highlight bound states of two $+1/2$ defects rotating around one another. The sub panels show four time snapshots of one of the bound states at $t=-4\Delta t$, $t=-2\Delta t$, $t=0$, and $t=2\Delta t$, where $t=0$ is the time in the main snapshot. (b) Three time snapshots of a confined continuum simulation with only two $+1/2$ defects with $\alpha/\ell_{h}^{2}=0.2$, $\ell_{h}=10$, $\varphi_{0}=\pi/8$, and domain radius $R=30$. (c) Scaled stable distance between defects $r_{0}/\ell_{h}$ in a continuum simulation for various values of $\alpha/\ell_{h}^{2}$ with $\ell_{h}=10$ and $\varphi_{0}=\pi/8$. The dashed line is a reproduction of the analytically derived curve in Fig. 3(b), $r_{0}(\varphi_{0}=\pi/8)/\ell_{h}+1.2$.
@@
By simulating a continuum system under circular confinement, we can perform a rough comparison to the scaling of $r_{0}$ with $\alpha$ shown in Fig. 3(b). In Fig. 4(b), we plot several time snapshots of a continuum simulation with $\alpha/\ell_{h}^{2}=0.2$, $\ell_{h}=10$, $\varphi_{0}=\pi/8$, and a confining radius of $R=30$. In the continuum simulations, two defects rotate around one another at a fixed distance. As predicted by the analytical equations, this distance depends on the activity. In Fig. 4(c) we plot the critical distance $r_{0}$ versus $\alpha/\ell_{h}^{2}$ for several continuum simulations. We also plot the analytically derived curve from Fig. 3(b), $r_{0}(\varphi_{0}=\pi/8)/\ell_{h}+1.2$, where the vertical shift in the curve is made to highlight the similarities and differences in the shape of the relationship. Both the continuum simulations and analytic model show a relatively rapid drop in $r_{0}$ for small $\alpha$; however, at larger activities $r_{0}$ falls off more rapidly with $\alpha$ in the continuum simulations than in the analytic model, which we attribute to nonlinear effects produced by other defects in close proximity to the defect pair in the continuum simulations. When the defects become close enough together that their cores begin to overlap, our approximations break down. In particular, in the continuum model, defects are able to merge to form higher order defects, as is visible in several of the boxed regions in Fig. 4(a). We surmise that the quantitative difference (the vertical shift between the analytically predicted behavior and the continuum simulations) arises due to effects of the boundaries on the continuum simulations. We use $\mathbf{u}=0$ boundary conditions for the flow field in the confined continuum simulations, but Eq. (22) corresponds to defects in an infinite, not bounded, domain. These boundary conditions reduce the flow magnitude and spatial extent, and, hence, increase the critical distance between defects. The vertical shift factor, here found to be $\approx 1.2$, likely depends on both $\ell_{h}$ and the size of the computational domain, though a comprehensive study of this effect is outside the scope of the present work. Additionally, because defects can spontaneously nucleate for higher activities in the continuum simulations, we cannot probe the high activity scaling result. In the lower activity regime, however, we obtain good qualitative agreement between our analytic model and the behavior of defects in continuum simulations and experiments.
@@ 3.2 Two $-1/2$ Defects
$\mathbf{v}^{--}=\left[\frac{1}{r}-\frac{\alpha}{16\ell_{h}^{2}}\chi_{1}\left(\frac{r}{2\ell_{h}}\right)\right]\mathbf{\hat{r}}\\ +\frac{1}{4\ell_{h}}\left[\beta_{3}\left(\frac{r}{2\ell_{h}}\right)\mathbf{\hat{c}}+\beta_{4}\left(\frac{r}{2\ell_{h}}\right)\mathbf{\hat{d}}\right]-\frac{\alpha}{16\ell_{h}^{2}}\chi_{4}\left(\frac{r}{2\ell_{h}}\right)\mathbf{\hat{f}}$ (25)
@@
where $\mathbf{\hat{c}}=\left[\cos(2\theta-2\varphi_{0}),-\sin(2\theta-2\varphi_{0})\right]$, $\mathbf{\hat{d}}=\left[\cos(4\theta-2\varphi_{0}),\sin(4\theta-2\varphi_{0})\right]$, and $\mathbf{\hat{f}}=\left[\cos(5\theta-4\varphi_{0}),-\sin(5\theta-4\varphi_{0}))\right]$. As with the two $+1/2$ defects, it is instructive to decompose the velocity into components parallel and transverse to the direction between the defects:
@@
$v_{||}^{-}=\frac{1}{r}-\frac{\alpha}{16\ell_{h}^{2}}\chi_{1}\\ +\frac{\alpha}{4\ell_{h}}\left(\beta_{3}+\beta_{4}\right)\cos(3\theta-2\varphi_{0})-\frac{\alpha}{16\ell_{h}^{2}}\chi_{3}\cos(6\theta-4\varphi_{0})$ (26)
@@
$v_{\perp}^{-}=-\frac{\alpha}{4\ell_{h}}\left(\beta_{3}-\beta_{4}\right)\sin(3\theta-2\varphi_{0})\\ +\frac{\alpha}{16\ell_{h}^{2}}\chi_{3}\sin(6\theta-4\varphi_{0}).$ (27)
@@
Fig.  5: (a) Scaled critical distance $r_{0}/\ell_{h}$ at which the velocity of $-1/2$ defects $\mathbf{v}=0$ for $03\theta-2\varphi_{0}=$ 0 versus $\alpha/\ell_{h}^{2}$. (b) $r_{0}/\ell_{h}$ versus $\alpha/\ell_{h}^{2}$ for two $-1/2$ defects in confined continuum simulations with $\ell_{h}=10$ and domain radius $R=30$. The dashed curve is a reproduction of the curve from (a), given by $r_{0}/\ell_{h}+2$. (c) Example trajectories of two $-1/2$ defects with initial $x=-0.5\ell_{h}$ and $x=0.5\ell_{h}$ at $\varphi_{0}=0.001$.
@@
Equation (27) shows that the transverse component of the velocity will be zero if $3\theta-2\varphi_{0}=n\pi$. In this case, two $-1/2$ defects can only move toward or away from one another. From Eq. (26), we find that if $3\theta-2\varphi_{0}=\pi+2n\pi$, the defects repel each other. Geometrically, this occurs when two $-1/2$ defects “point” towards one another if viewed as having a triangular shape. On the other hand, when the defects point away from one another for $3\theta-2\varphi=2n\pi$, the hydrodynamic interactions compete with the Coulomb force on the defects and there is a critical distance $r_{0}$ at which $v_{||}=0$. In Fig. 5(a) we plot this critical distance as a function of $\alpha/\ell_{h}^{2}$ for $\ell_{h}=10$. For two $-1/2$ defects, $r_{0}$ rapidly decreases as $\alpha/\ell_{h}^{2}$ increases until $\alpha/\ell_{h}\approx 1$ and the value of $r_{0}$ saturates. As we did with the $+1/2$ defects, in Fig. 5(b) we plot $r_{0}/\ell_{h}$ for several confined continuum simulations with two $-1/2$ defects. The dashed line in Fig. 5(b) is the curve from Fig. 5(a) with a vertical shift, $r_{0}/\ell_{h}+2$, showing that apart from the vertical shift, the finite system continuum simulations behave qualitatively similarly to the curve derived analytically for an infinite domain. As with the $+1/2$ defects in the previous section, we suspect the quantitative difference to result from the confinement of the continuum simulations, where the flow speeds are reduced even more than in the case of two $+1/2$ defects. We expect that the value of the vertical shift, found to be $2$ for the parameters used here, should be dependent on $\ell_{h}$ and the domain size of the confined system.
@@
Although a critical distance between $-1/2$ defects exists, the velocity equations do not support stable, bound defect pairs. To see this, note that if the configuration is rotated by $-\theta$, $\varphi_{0}\to\varphi_{0}+(1/2)\theta$ due to the $-1/2$ winding character of the defect. The angular argument in Eqs. (26) and (27) is then transformed as $3\theta-2\varphi_{0}\to-\theta-2\varphi_{0}$. This means that transverse movements between defects will change the angular argument. Further, expanding Eq. (27) for small $3\theta-2\varphi_{0}$ shows that transverse velocities will increase for $3\theta-2\varphi_{0}\neq 0$. We illustrate this instability in Fig. 5(c) by plotting the computed trajectories of two $-1/2$ defects with initial positions $(-0.5\ell_{h},0)$ and $(0.5\ell_{h},0)$, for $\varphi_{0}=0.001$, $\ell_{h}=10$ and $\alpha/\ell_{h}^{2}=0.5$. Initially the defects move towards each other until they rotate enough for the interaction to become repulsive. They then repel each other along a line such that $3\theta-2\varphi_{0}=\pi$. This instability provides an explanation for why bound states of $-1/2$ defect pairs are not observed in continuum simulations and experiments.
@@
Unlike the previous two cases we have examined, ±$\pm 1/2$ defect pairs interact non-reciprocally. That is, $\mathbf{v}^{+-}\neq-\mathbf{v}^{-+}$ and, thus, the predicted trajectories will not be symmetric. The velocities of the defects are given by
$\mathbf{v}^{+-}=-\frac{1}{r}\mathbf{\hat{r}}-\frac{\alpha}{16\ell_{h}^{2}}\chi_{1}\left(\frac{r}{2\ell_{h}}\right)\mathbf{\hat{r}}^{*}-\frac{\pi\alpha}{4\ell_{h}}\mathbf{\hat{p}}\\ +\frac{\alpha}{4\ell_{h}}\left[\beta_{3}(\left(\frac{r}{2\ell_{h}}\right)\mathbf{\hat{c}}+\beta_{4}\left(\frac{r}{2\ell_{h}}\right)\mathbf{\hat{d}}\right]-\frac{\alpha}{16\ell_{h}^{2}}\chi_{3}\left(\frac{r}{2\ell_{h}}\right)\mathbf{\hat{f}},$ (28)
$\mathbf{v}^{-+}=-\frac{1}{r}\mathbf{\hat{r}}-\frac{\alpha}{16\ell_{h}^{2}}\chi_{1}\left(\frac{r}{2\ell_{h}}\right)\mathbf{\hat{r}}^{*}\\ +\frac{\alpha}{4\ell_{h}}\left[\beta_{1}\left(\frac{r}{2\ell_{h}}\right)\mathbf{\hat{p}}+\beta_{2}\left(\frac{r}{2\ell_{h}}\right)\mathbf{\hat{a}}\right]-\frac{\alpha}{16\ell_{h}^{2}}\chi_{2}\left(\frac{r}{2\ell_{h}}\right)\mathbf{\hat{b}}$ (29)
where $\mathbf{\hat{r}}^{*}=(\cos\theta,-\sin\theta)$ and we note that we have written the velocities in coordinates where the origin is located at the other defect.
We first analyze the interaction by examining the limiting case where the global phase is $\varphi_{0}=\pi/2$ and the axes are oriented so that $\theta=0$. In this case the comet head of the $+1/2$ defect is pointing “away” from the $-1/2$ defect [see Fig. 6(a)]. The velocity difference, and, hence, the rate at which the distance between defects changes, is
$\frac{dr}{dt}=-\frac{2}{r}+\frac{\alpha}{4\ell_{h}}\left(\pi-\beta_{1}+\beta_{2}-\beta_{3}-\beta_{4}\right)-\frac{\alpha}{16\ell_{h}^{2}}\left(\chi_{2}+\chi_{3}\right).$ (30)

### v2

@@ 3 Interaction between two active nematic defects
Here we analyze the interaction between active nematic defects with the simplifying assumption that only two defects exist in an infinite domain. In each case we set our coordinates so that the defect of interest is at polar coordinates $(r,\theta)$ and the other defect is at $(0,0)$. We will also assume “optimal” orientation between defects 9. This set of orientations minimizes the elastic free energy of the configuration. Further, given a global phase $\varphi_{0}$, the orientation of each defect may be determined by the defect positions:
@@ 3.1 Two $+1/2$ Defects
where $\mathbf{\hat{r}}$ is the unit vector pointing to the defect of interest from the other defect, $\mathbf{\hat{a}}=\left[\cos(\theta-2\varphi_{0}),\sin(\theta-2\varphi_{0})\right]$, and $\mathbf{\hat{b}}=\left[\cos(\theta-4\varphi_{0}),\sin(\theta-4\varphi_{0})\right]$. Components of Eq. (22) proportional to $\beta_{i}$ result from advection while components proportional to $\chi_{i}$ result from the interaction with shear flows. We note that the other defect will have exactly the opposite velocity of the defect under consideration.
@@
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}v_{||}^{+}=\frac{1}{r}-\frac{\alpha}{16\ell_{h}^{2}}\chi_{1}}\\ {\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}+\frac{\alpha}{4\ell_{h}}\left(-\pi+\beta_{1}+\beta_{2}\right)\cos(2\varphi_{0})-\frac{\alpha}{16\ell_{h}^{2}}\chi_{2}\cos(4\varphi_{0})}$ (23)
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}v_{\perp}^{+}=\frac{\alpha}{4\ell_{h}}\left(-\pi+\beta_{1}-\beta_{2}\right)\sin(2\varphi_{0})+\frac{\alpha}{16\ell_{h}^{2}}\chi_{2}\sin(4\varphi_{0})}$ (24)
where we have suppressed the arguments of $\beta_{1},\beta_{2},\chi_{1},\chi_{2}$ for brevity. Equations (23) and (24) are independent of $\theta$, the angle between defects, and the angular dependence is given by the fixed global phase $\varphi_{0}$. Thus, since $v_{\perp}$ is proportional to the change of angle between defects, the model predicts that defects will rotate around one another as long as $v_{\perp}\neq 0$. We note that the independence of $\theta$ in the defect dynamics is a consequence of the unique rotational symmetry of the far-field director for total topological charge $+1$. As we will show in later sections, other defect interactions do not share this property.
@@
The behavior of $r_{0}$ may be attributed primarily to the self-advection part of Eq. (23), which is the dominant term in the equation. It does not depend on the distance between defects, and it contributes most strongly at high activities and smaller $\varphi_{0}$. However, we note that the shear flow interaction contributes most strongly to the behavior near $\varphi_{0}=\pi/4$ where the advection terms do not contribute to Eq. (23). If $\lambda=0$, $r_{0}\to\infty$ for all values of $\alpha$ at $\varphi_{0}=\pi/4$.
@@
The effective attraction between positive defects caused by hydrodynamic effects can lead to dynamical states in which $+1/2$ defects continuously rotate around one another, or even merge to form a stable higher order $+1$ defect, similar to behaviors predicted with a different analytical model in Ref. 49. While our calculations above predict this attraction for the simplest case of a system with only two defects, this behavior should, in principle, occur even when other defects are present in the system if the two defects are far from the other defects. Indeed, bound states of two $+1/2$ defects may be observed in continuum simulations at high activities, as illustrated in Fig. 4(a) where we show a time snapshot of the scalar order parameter and director field in a system with periodic boundary conditions, $\alpha/\ell_{h}^{2}=2$, and $\ell_{h}=10$ (for details of the continuum simulations, see the Appendix). We have marked the bound states of $+1/2$ defects in green boxes in Fig. 4(a), and for one of these boxes the subpanels provide several time snapshots illustrating two bound $+1/2$ defects rotating around one another and eventually merging. We note that the bound states are short-lived in the continuum simulation due to interactions with other defects. However, it has been shown in numerical simulations that such bound states may be longer lived and ordered if subjected to confinement 59, 60, 61. Partially stable bound $+1/2$ defect pairs have also been observed experimentally under circular confinement 62 and in recently discovered acoustically actuated active nematics with no confinement, where the activity may be much larger than in biologically based active nematics 63.
@@
Fig.  4: (a) Time snapshot of an active nematic continuum simulation with periodic boundary conditions, $\alpha/\ell_{h}^{2}=2$, and $\ell_{h}=10$. The color represents the strength of the scalar order parameter $S$ and the white lines indicate the local director. The green boxes highlight bound states of two $+1/2$ defects rotating around one another. The sub panels show four time snapshots of one of the bound states at $t=-4\Delta t$, $t=-2\Delta t$, $t=0$, and $t=2\Delta t$, where $t=0$ is the time in the main snapshot. The circles indicate the $+1/2$ defects that rotate around one another. (b) Three time snapshots of a confined continuum simulation with only two $+1/2$ defects with $\alpha/\ell_{h}^{2}=0.2$, $\ell_{h}=10$, $\varphi_{0}=\pi/8$, and domain radius $R=30$. (c) Scaled stable distance between defects $r_{0}/\ell_{h}$ in a continuum simulation for various values of $\alpha/\ell_{h}^{2}$ with $\ell_{h}=10$ and $\varphi_{0}=\pi/8$ (orange points). The dashed line is a reproduction of the analytically derived curve from Eq. (23) (solid line), with a vertical shift of $1.2$.
@@
By simulating a continuum system under circular confinement, we can perform a rough comparison to the scaling of $r_{0}$ with $\alpha$ shown in Fig. 3(b). In Fig. 4(b), we plot several time snapshots of a continuum simulation with $\alpha/\ell_{h}^{2}=0.2$, $\ell_{h}=10$, $\varphi_{0}=\pi/8$, and a confining radius of $R=30$. In the continuum simulations, two defects rotate around one another at a fixed distance. As predicted by the analytical equations, this distance depends on the activity. In Fig. 4(c) we plot the critical distance $r_{0}$ versus $\alpha/\ell_{h}^{2}$ for several continuum simulations. We also plot the analytically derived curve from Fig. 3(b), $r_{0}(\varphi_{0}=\pi/8)/\ell_{h}+1.2$, where the vertical shift in the curve is made to highlight the similarities and differences in the shape of the relationship. Both the continuum simulations and analytic model show a relatively rapid drop in $r_{0}$ for small $\alpha$; however, at larger activities $r_{0}$ falls off more rapidly with $\alpha$ in the continuum simulations than in the analytic model, which we attribute to nonlinear effects produced by the close proximity of the defect pair in the continuum simulations. When the defects become close enough together that their cores begin to overlap, our approximations break down. In particular, in the continuum model, defects are able to merge to form higher order defects, as is visible in several of the boxed regions in Fig. 4(a). We surmise that the quantitative difference (the vertical shift between the analytically predicted behavior and the continuum simulations) arises due to effects of the boundaries on the continuum simulations. We use $\mathbf{u}=0$ boundary conditions for the flow field in the confined continuum simulations, but Eq. (22) corresponds to defects in an infinite, not bounded, domain. These boundary conditions reduce the flow magnitude and spatial extent, and, hence, increase the critical distance between defects. The vertical shift factor, here found to be $\approx 1.2$, likely depends on both $\ell_{h}$ and the size of the computational domain, though a comprehensive study of this effect is outside the scope of the present work. Additionally, because defects can spontaneously nucleate for higher activities in the continuum simulations, we cannot probe the high activity scaling result. In the lower activity regime, however, we obtain good qualitative agreement between our analytic model and the behavior of defects in continuum simulations and experiments.
@@ 3.2 Two $-1/2$ Defects
$\mathbf{v}^{--}=\left[\frac{1}{r}-\frac{\alpha}{16\ell_{h}^{2}}\chi_{1}\left(\frac{r}{2\ell_{h}}\right)\right]\mathbf{\hat{r}}\\ +\frac{\alpha}{4\ell_{h}}\left[\beta_{3}\left(\frac{r}{2\ell_{h}}\right)\mathbf{\hat{c}}+\beta_{4}\left(\frac{r}{2\ell_{h}}\right)\mathbf{\hat{d}}\right]-\frac{\alpha}{16\ell_{h}^{2}}\chi_{3}\left(\frac{r}{2\ell_{h}}\right)\mathbf{\hat{f}}$ (25)
@@
where $\mathbf{\hat{c}}=\left[\cos(3\theta-2\varphi_{0}),-\sin(3\theta-2\varphi_{0})\right]$, $\mathbf{\hat{d}}=\left[\cos(5\theta-2\varphi_{0}),\sin(5\theta-2\varphi_{0})\right]$, and $\mathbf{\hat{f}}=\left[\cos(7\theta-4\varphi_{0}),-\sin(7\theta-4\varphi_{0})\right]$. As with the two $+1/2$ defects, it is instructive to decompose the velocity into components parallel and transverse to the direction between the defects:
@@
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}v_{||}^{-}=\frac{1}{r}-\frac{\alpha}{16\ell_{h}^{2}}\chi_{1}}\\ {\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}+\frac{\alpha}{4\ell_{h}}\left(\beta_{3}+\beta_{4}\right)\cos(4\theta-2\varphi_{0})-\frac{\alpha}{16\ell_{h}^{2}}\chi_{3}\cos(8\theta-4\varphi_{0})}$ (26)
@@
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}v_{\perp}^{-}=-\frac{\alpha}{4\ell_{h}}\left(\beta_{3}-\beta_{4}\right)\sin(4\theta-2\varphi_{0})}\\ {\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}+\frac{\alpha}{16\ell_{h}^{2}}\chi_{3}\sin(8\theta-4\varphi_{0}).}$ (27)
@@
Fig.  5: (a) Scaled critical distance $r_{0}/\ell_{h}$ at which the velocity of $-1/2$ defects $\mathbf{v}=0$ for $04\theta-2\varphi_{0}=$ 0 versus $\alpha/\ell_{h}^{2}$. The solid line shows the calculated $r_{0}/\ell_{h}$ from Eq. (26). Orange points show $r_{0}/\ell_{h}$ versus $\alpha/\ell_{h}^{2}$ for two $-1/2$ defects in confined continuum simulations with $\ell_{h}=10$ and domain radius $R=30$. The dashed curve is a reproduction of the solid curve with a vertical shift of $2$. (b) Example trajectories of two $-1/2$ defects with initial $x=-0.5\ell_{h}$ and $x=0.5\ell_{h}$ at $\varphi_{0}=0.001$.
@@
Equation (27) shows that the transverse component of the velocity will be zero if $4\theta-2\varphi_{0}=n\pi$. In this case, two $-1/2$ defects can only move toward or away from one another. From Eq. (26), we find that if $4\theta-2\varphi_{0}=(2n+1)\pi$, the defects repel each other. Geometrically, this occurs when two $-1/2$ defects “point” towards one another if viewed as having a triangular shape. On the other hand, when the defects point away from one another for $4\theta-2\varphi=2n\pi$, the hydrodynamic interactions compete with the Coulomb force on the defects and there is a critical distance $r_{0}$ at which $v_{||}=0$. In Fig. 5(a) we plot this critical distance as a function of $\alpha/\ell_{h}^{2}$ for $\ell_{h}=10$. For two $-1/2$ defects, $r_{0}$ rapidly decreases as $\alpha/\ell_{h}^{2}$ increases until $\alpha/\ell_{h}\approx 1$ and the value of $r_{0}$ saturates. As we did with the $+1/2$ defects, in Fig. 5(a) we plot $r_{0}/\ell_{h}$ for several confined continuum simulations with two $-1/2$ defects. The dashed line in Fig. 5(a) is the solid curve in the figure with a vertical shift, $r_{0}/\ell_{h}+2$, showing that apart from the vertical shift, the finite system continuum simulations behave qualitatively similarly to the curve derived analytically for an infinite domain. As with the $+1/2$ defects in the previous section, we suspect the quantitative difference to result from the confinement of the continuum simulations, where the flow speeds are reduced even more than in the case of two $+1/2$ defects. We expect that the value of the vertical shift, found to be $2$ for the parameters used here, should be dependent on $\ell_{h}$ and the domain size of the confined system.
@@
Although a critical distance between $-1/2$ defects exists, the velocity equations do not support stable, bound defect pairs. This is because the dynamics are explicitly dependent on $\theta$, and, in particular, $v_{\perp}$ changes as the defects rotate around one another. Further, expanding around the fixed points of Eq. (27) shows that $4\theta-2\varphi_{0}=2n\pi$ are unstable fixed points, while $4\theta-2\varphi_{0}=(2n+1)\pi$ are stable fixed points. The model thus predicts that active $-1/2$ defects will tend to reorient toward a stable relative orientation when interacting. We illustrate this instability in Fig. 5(b) by plotting the computed trajectories of two $-1/2$ defects with initial positions $(-0.5\ell_{h},0)$ and $(0.5\ell_{h},0)$, for $\varphi_{0}=0.001$, $\ell_{h}=10$ and $\alpha/\ell_{h}^{2}=0.5$. Initially the defects move towards each other until they rotate enough for the interaction to become repulsive. They then repel each other along a line such that asymptotically approaches $4\theta-2\varphi_{0}=\pi$. This instability provides an explanation for why bound states of $-1/2$ defect pairs are not observed in continuum simulations and experiments.
@@
Unlike the previous two cases we have examined, ±$\pm 1/2$ defect pairs interact non-reciprocally 12. That is, $\mathbf{v}^{+-}\neq-\mathbf{v}^{-+}$ and, thus, the predicted trajectories will not be symmetric. The velocities of the defects are given by
$\mathbf{v}^{+-}=\left[-\frac{1}{r}-\frac{\alpha}{16\ell_{h}^{2}}\chi_{1}\left(\frac{r}{2\ell_{h}}\right)\right]\mathbf{\hat{r}}-\frac{\pi\alpha}{4\ell_{h}}\mathbf{\hat{p}}\\ +\frac{\alpha}{4\ell_{h}}\left[\beta_{3}\left(\frac{r}{2\ell_{h}}\right)\mathbf{\hat{a}^{*}}+\beta_{4}\left(\frac{r}{2\ell_{h}}\right)\mathbf{\hat{c}^{*}}\right]-\frac{\alpha}{16\ell_{h}^{2}}\chi_{3}\left(\frac{r}{2\ell_{h}}\right)\mathbf{\hat{g}},$ (28)
$\mathbf{v}^{-+}=\left[-\frac{1}{r}-\frac{\alpha}{16\ell_{h}^{2}}\chi_{1}\left(\frac{r}{2\ell_{h}}\right)\right]\mathbf{\hat{r}}\\ +\frac{\alpha}{4\ell_{h}}\left[\beta_{1}\left(\frac{r}{2\ell_{h}}\right)\mathbf{\hat{a}^{*}}+\beta_{2}\left(\frac{r}{2\ell_{h}}\right)\mathbf{\hat{c}^{*}}\right]-\frac{\alpha}{16\ell_{h}^{2}}\chi_{2}\left(\frac{r}{2\ell_{h}}\right)\mathbf{\hat{h}}$ (29)
where $\mathbf{\hat{a}^{*}}=[\cos(\theta-2\varphi_{0}),-\sin(\theta-2\varphi_{0})]$, $\mathbf{\hat{c}^{*}}=[\cos(3\theta-2\varphi_{0}),\sin(3\theta-2\varphi_{0})]$, $\mathbf{\hat{g}}=[\cos(5\theta-4\varphi_{0}),\sin(5\theta-4\varphi_{0})]$, $\mathbf{\hat{h}}=[\cos(3\theta-4\varphi_{0}),-\sin(3\theta-4\varphi_{0})]$, and we note that we have written the velocities in coordinates where the origin is located at the other defect.
We first analyze the interaction by examining the limiting case where the global phase is $\varphi_{0}=\pi/2$ and the axes are oriented so that $\theta=0$ for the $+1/2$ defect and $\theta=\pi$ for the $-1/2$ defect. In this case the comet head of the $+1/2$ defect is pointing “away” from the $-1/2$ defect [see Fig. 6(a)]. The velocity difference, and, hence, the rate at which the distance between defects changes, is
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\frac{dr}{dt}=-\frac{2}{r}+\frac{\alpha}{4\ell_{h}}\left(\pi-\beta_{1}+\beta_{4}\right)-\frac{\alpha}{16\ell_{h}^{2}}\left(2\chi_{1}+\chi_{2}+\chi_{3}\right).}$ (30)